Given angle H and angle J are complementary angles and the measure of angle H is five less than four times the measure of angle J what is the angle of each measure?
Answer by cleomenius(959)

x + 4x - 5 = 90
5x - 5 = 90
5x = 95
x = 19 degrees. This is angle H.
4(19) = 76 - 5 = 71 Degrees.
19 + 71 = 90 degrees.
This does check.
Cleomenius.